OpenStudy (anonymous):

Write the equation of a line in slope-intercept form that is perpendicular toy=-5/6x+2 and passes through (12, -4).Answer

OpenStudy (aroub):

y=-5/6x+2 y-y1=m(x-x1) , m= 6/5 , (12,-4) y+4=6/5(x-12)

OpenStudy (anonymous):

in slope intercept form its, x/(31/3)-y/(5/92)=1